OpenStudy (anonymous):

Find |u + v|, |u − v|,and |u| − |v|. u = −2i + 4j, v = i − 5j

OpenStudy (phi):

can you first figure out what u+v is ?

OpenStudy (anonymous):

<-1,-1>

OpenStudy (phi):

now use pythagoras to find the lenghth

OpenStudy (phi):

I get |u+v| = sqrt(2)

OpenStudy (anonymous):

that what i got but i dont know what to do next

OpenStudy (phi):

you do |u-v|

OpenStudy (phi):

what is u-v?

OpenStudy (anonymous):

\[\sqrt{10}\]

OpenStudy (phi):

u = −2i + 4j, v = i − 5j -2i -i + 4j - -5j -3i +9j

OpenStudy (phi):

now find | -3i + 9j| = sqrt(9+81)= sqrt(90)

OpenStudy (phi):

or 3 sqrt(10)

OpenStudy (phi):

how did you get sqrt(10) ?

OpenStudy (anonymous):

i did 4+-5 and and got -1 for b

OpenStudy (anonymous):

but i see what i did wrong

OpenStudy (phi):

now the last problem first find | u | then |v|

OpenStudy (anonymous):

u =sqrt 20,v=sqrt26

OpenStudy (anonymous):

2sqrt5-sqrt26

OpenStudy (phi):

that's what I got. an ugly answer, but it is what it is...

OpenStudy (anonymous):

thank you for the help it help alot